# Simultaneous extension of the polynomial Central Sets Theorem and the polynomial Milliken-Taylor theorem

### v1

@@ Abstract
In this article we find the polynomial extension of Beigelböck’s [1] multidimensional Central Sets Theorem (which was a common extension of Furstenberg’s Central Sets Theorem and Milliken-Taylor theorem). Then we derive a common extension of Bergelson, Hindman, and William’s [2] polynomial Milliken-Taylor theorem, and the polynomial multidimensional Central Sets Theorem. We also find a sufficient condition over two distinct multidimensional polynomials for which their Milliken-Taylor image lies in the same color. This result gives a partial polynomial extension of Deuber, Hindman, Leader, and Lefmann’s [5] result on separating the Milliken-Taylor system.
@@
Keywords: Central sets, Central Sets Theorem, Milliken-Taylor system, Algebra of the Stone-Čech Compactifications of discrete semigroups.
@@ 1   Introduction
Throughout our article by $\mathbb{N}$, we denote the set of all positive integers. For any $r\in\mathbb{N},$ a $r$-coloring of a set $X$ is a partition of $X$ into $r$ disjoint sets. By finite coloring we mean any $r$-coloring, for any $r\in\mathbb{N}$. Let $(S,\cdot)$ be a semigroup, and $\mathcal{F}$ be a family of subsets of $S.$ We say $\mathcal{F}$ is partition regular (P.R.) if, for any finite coloring of $S,$ there exists a member of $F\in\mathcal{F}$ in the same color, i.e. $F$ is monochromatic. A cornerstone result in arithmetic Ramsey theory is the van der Waerden’s theorem [21], which states that the set of all arithmetic progression of a given length is P.R. Another important theorem was due to Schur [19] which states that the collection $\{\{x,y,x+y\}:x\neq y\in\mathbb{Z}\}$ is P.R. A nonlinear version of van der Waerden’s theorem, known as the Polynomial van der Waerden’s Theorem [^3] Actually, the authors proved a generalized version of the much stronger Szémeredi’s Theorem, but we are not going to discuss it in this paper. was established by Bergelson and Liebman in [3], using the methods of topological dynamics and PET induction. Throughout this article, we assume $\mathbb{P}$ be the set of all polynomials with no constant term.
@@
Let $r\in\mathbb{N},$ and $\mathbb{N}=\bigcup_{i=1}^{r}C_{i}$ be a $r$-coloring of $\mathbb{N}$. Then for any finite collection of polynomials $F$ in $\mathbb{P}$, there exist $a,d\in\mathbb{N}$ and $1\leq j\leq r$ such that $\{a+p(d):p\in F\}\subset C_{j}.$
An infinitary extension of Schur’s theorem inspires the following definition of IP sets. For any set $X$, let $\mathcal{P}_{f}\left(X\right)$ be the collection of all nonempty finite subsets of $X.$
Definition 1.2 (IP Sets).
Let $(S,+)$ be a commutative semigroup. A set $A\subseteq S$ is said to be an $IP$ set if there exists an injective sequence $\langle x_{n}\rangle_{n\in\mathbb{N}}$ in $S$ such that
$A=FS\left(\langle x_{n}\rangle_{n\in\mathbb{N}}\right)=\left\{\sum_{n\in\alpha}x_{n}:\alpha\in\mathcal{P}_{f}\left(\mathbb{N}\right)\right\}.$
Abbreviately for $\alpha\in\mathcal{P}_{f}\left(\mathbb{N}\right),$ we write $x_{\alpha}=\sum_{n\in\alpha}x_{n}$.
The following theorem is a cornerstone theorem in Ramsey theory, often known as the Hindman theorem [9].
Theorem 1.3 (Hindman Theorem).
@@
For details on the classical approach to the Ramsey theory we refer to the book [8]. In our article we use Ultrafilter theory. So, before we proceed, we need to recall some basic results from the Stone-Čech compactification of discrete semigroups, sometimes known as Topological algebra/ theory of ultrafilter.
Topological algebra
Ultrafilters are set-theoretic objects which are intimately related to the Ramsey theory. In this section, we give a brief introduction to this theory. For details we refer to the book [13] to the readers.
A filter $\mathcal{F}$ over any nonempty set $X$ is a collection of subsets of $X$ such that
@@
  • if $X=\cup_{i=1}^{r}A_{i}$ is any finite partition of $X$, then there exists $i\in\{1,2,\ldots,r\}$ such that $A_{i}\in p.$
Let $(S,\cdot)$ be any discrete semigroup. Let $\beta S$ be the collection of all ultrafilters. For every $A\subseteq S,$ define $\overline{A}=\{p\in\beta S:A\in p\}.$ Now one can check that the collection $\{\overline{A}:A\subseteq S\}$ forms a basis for a topology. This basis generates a topology over $\beta S.$ We can extend the operation ``"``"`$\cdot$"` ` ⋅ " of $S$ over $\beta S$ as: for any $p,q\in\beta S,$ $A\in p\cdot q$ if and only if $\{x:x^{-1}A\in q\}\in p.$ With this operation ``"``"`$\cdot$"` ` ⋅ ", $(\beta S,\cdot)$ becomes a compact Hausdorff right topological semigroup. One can show that $\beta S$ is nothing but the Stone-Čech compactification of $S.$ Hence Ellis’s theorem guarantees that there exist idempotents in $(\beta S,\cdot)$. The set of all idempotents in $(\beta S,\cdot)$ is denoted by $E\left((\beta S,\cdot)\right).$ It can be shown that every member of the idempotents of $(\beta S,\cdot)$ contains an $IP$ set, which means every idempotent witnesses Hindman’s theorem. Using Zorn’s lemma one can show that $(\beta S,\cdot)$ contains minimal left ideals (minimal w.r.t. the inclusion). A well-known fact is that the union of such minimal left ideals is a minimal two-sided ideal, denoted by $K(\beta S,\cdot).$ Here we recall a few well-known classes of sets that are relevant to our work.
@@
  • $A$ is a thick set if for any finite subset $F\subset S$, there exists an element $x\in S$ such that $Fx=\{fx:f\in F\}\subset A$;
  • $A$ is a syndetic set if there exists a finite set $F\subset S$ such that $S=\bigcup_{x\in F}x^{-1}A$, where $x^{-1}A=\{y:xy\in A\}$;
  • $A$ is piecewise syndetic set if there exists a finite set $F\subset S$ such that $\bigcup_{x\in F}x^{-1}A$ is a thick set. It is well known that $A$ is piecewise syndetic if and only if there exists $p\in K(\beta S,\cdot)$ such that $A\in p.$
  • $A$ is central set if it belongs to a minimal idempotent in $\beta S$.
It can be proved that a set $A$ is thick if and only if there exists a left ideal $L$ such that $L\subseteq\overline{A}$. And a set $A$ is syndetic if and only if for every left ideal $L$, $L\cap\overline{A}\neq\emptyset.$ Note that if $f:(S,\cdot)\rightarrow(T,\cdot)$ be any map between discrete semigroups. Then $\tilde{f}:\beta S\rightarrow\beta T$ be the continuous extension of $f$ defined by $\tilde{f}(p)=\tilde{f}(p-\lim x)=p-\lim f(x),$ where $p-\lim f(x)=y$ if and only if for every neighbourhood $U$ of $y$, $\{x:f(x)\in U\}\in p.$
@@
The tensor product of ultrafilters was introduced in [15]. Let $(S,\cdot)$ and $(T,\cdot)$ be two discrete semigroups. In this section, we recall materials from [14].
@@
We can also extend the above definition of tensor products for multiple ultrafilters. The following definition uses inductive arguments.
@@
In a similar fashion we can say $\bigotimes_{i=1}^{k}p_{i}=\lim_{s_{1}\rightarrow p_{1}}\cdots\lim_{s_{k}\rightarrow p}(s_{1},\ldots,s_{k}).$
In the literature, the tensor product trick has been used to prove multi-dimensional Ramsey theoretic results including Milliken-Taylor theorem, and it’s polynomial extensions. For details, one can see [2, 13, 16]. In [1], Beigelböck proved a simultaneous extension of the Central Sets Theorem and Milliken-Taylor theorem using a little different approach described in [2, Chapter 18]. One of our main motivations is to study the tensor product of minimal idempotent ultrafilters, which was not extensively covered earlier. As a consequence we are able to find polynomial extensions of several multidimensional results in Ramsey theory. Before we proceed, we need the following technical Lemma, which is essential for our purpose.
@@
[2, Lemma 4.3] Let $(S,\cdot)$ be a semigroup, let $m\in\mathbb{N}$, and let $p\in\beta S.$ Let $A\in\bigotimes_{j=1}^{m}p.$ Then for $j\in\{1,2,\ldots,m\}$ there exists $D_{j}:S^{j-1}\rightarrow\mathcal{P}(S)$ such that
@@
    for $j\in\{1,2,\ldots,m\},$ if for each $s\in\{1,2,\ldots,j-1\},$ $w_{s}\in D_{s}(w_{1},\ldots,w_{s-1}),$ then $D_{j}(w_{1},\ldots,w_{j-1})\in p;$
@@
Central Sets Theorem and its polynomial extensions
Before we proceed to our results, we need to recall the Central Sets Theorem. After the foundation of both van der Waerden’s and Hindman’s theorem, an immediate question appears if one can find a joint extension of both of these theorems. In [6], using the methods of Topological dynamics, Furstenberg defined the notions of Central Sets and proved that if $\mathbb{N}$ is finitely colored, then one of the color classes is Central. Then he proved a joint extension of both the van der Waerden’s and Hindman’s theorem, known as the Central Sets Theorem.
@@
Let $l\in\mathbb{N}$, and $A\subseteq\mathbb{N}$ be a central set. For each $i\in\{1,2,\ldots,l\}$ let $\langle x_{i,m}\rangle_{m=1}^{\infty}$ be a sequence in $\mathbb{\mathbb{Z}}$. Then there exist sequences $\langle b_{m}\rangle_{m=1}^{\infty}$ in $\mathbb{N}$ and $\langle K_{m}\rangle_{m=1}^{\infty}$ in $\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
@@
Later the above theorem was generalized in several directions. For details about the developments and their combinatorial applications, we refer to the book [13]. Before we state the stronger version of Theorem 1.8, we need the following definitions from [4].
@@
    $A$ is called a $J$-set if and only if for every $H\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, there exists $a\in\mathbb{N}$ and $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that for all $f\in H,$ $a+\sum_{t\in\beta}f(t)\in A.$
@@
It can be proved that $J$ sets are partition regular, and so $\mathcal{J}\neq\emptyset$. Again it is easy to verify that $(\mathcal{J},+)$ is a two sided ideal of $(\beta\mathbb{N},+)$, hence contains $K(\beta\mathbb{N},+).$ A set will be called $C$-set if it is a member of any idempotent in $(\mathcal{J},+)$. The following stronger version of the Central Sets Theorem was due to De, Hindman, and Strauss [4].
@@
Let $\tau=^{\mathbb{N}}\mathbb{N}$, and $C\subseteq\mathbb{N}$ be $C$ set. Then there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N}$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
@@
    $\sum_{i=1}^{m}\bigl{(}\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\bigr{)}\in C.$
In fact it was also shown in [4] that if a set $A\subset\mathbb{N}$ contains the conclusion of Theorem 1.10, then it is a $C$ set.
Polynomial extension
In [11, 12] Hindman and McCutcheon extensively studied the algebraic structure of VIP systems which are the generalizations of usual polynomials. They proved several results including the polynomial extension of Theorem 1.8. This was the joint extension of the polynomial van der Waerden’s theorem and the central sets theorem.
@@
Let $F\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, let $T\in\mathcal{P}_{f}(\mathbb{P})$ and let $A$ be a central subset of $\mathbb{N}$. Then there exist sequences $\langle b_{n}\rangle_{n=1}^{\infty}$ in $\mathbb{N}$ and $\langle H_{n}\rangle_{n=1}^{\infty}$ in $\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
@@
Later in [7], both of the authors of this paper and Luperi Bagilini found the polynomial extension of Theorem 1.10. To state this result we need the following definition.
@@
    $A$ is called a $J_{p}$-set if and only if for every $F\in\mathcal{P}_{f}(\mathbb{P})$, and every $H\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, there exists $a\in\mathbb{N}$ and $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that for all $P\in F$ and all $f\in H,$ $a+P\left(\sum_{t\in\beta}f(t)\right)\in A$
@@
    if $n\in\mathbb{N}$, $G_{1},G_{2},\cdots,G_{n}\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, $G_{1}\subsetneqq G_{2}\subsetneqq\cdots\subsetneqq G_{n}$ and for all $i\in\{1,2,\ldots,n\}$, $f_{i}\in G_{i}$, then for all $P\in T$,
@@
However likewise in Theorem 1.10, the converse direction of Theorem 1.13 is not known. By observing that any $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ is a subset of $\{1,\dots,n\}$ for some large enough $n\in\mathbb{N}$, we deduce the following seemingly stronger version of Theorem 1.13.
@@
Milliken-Taylor Theorem and its extensions
In $1975$-$76$, Milliken and Taylor independently proved a joint extension of both Hindman theorem, and Ramsey’s theorem. To state the Milliken-Taylor Theorem, we need to introduce some notation. Given $F,G\in\mathcal{P}_{f}(\mathbb{N})$, we write $F<G$ to mean that $\max F<\min G.$ When we say that a sequence $\langle H_{n}\rangle_{n}$ in $\mathcal{P}_{f}(\mathbb{N})$ is increasing if for every $n\in\mathbb{N}$, $\max H_{n}<\min H_{n+1}.$ We present the different terminology because these special cases arise frequently.
@@
    $[FS(\langle x_{n}\rangle_{n})]_{<}^{k}=\left\{\left(\sum_{t\in F_{1}}x_{t},\sum_{t\in F_{2}}x_{t},\ldots,\sum_{t\in F_{k}}x_{t}\right):F_{1}<F_{2}<\cdots<F_{k}\right\}.$
@@
    In a semigroup $(S,\cdot),$ $FP(\langle y_{n}\rangle_{n})$ is a product subsystem of $FP(\langle x_{n}\rangle_{n})$ if and only if there exists an increasing sequence $\langle H_{n}\rangle_{n}$ in $\mathcal{P}_{f}(\mathbb{N})$ such that for each $n\in\mathbb{N},$ $y_{n}=\Pi_{t\in H_{n}}x_{t}$ where the products $\Pi_{t\in H_{n}}x_{t}$ are computed in increasing order of indices. If the semigroup is commutative, we denote it by $(S,\cdot)$ and the product subsystem $FP(\langle y_{n}\rangle_{n})$ is denoted by $FS(\langle y_{n}\rangle_{n})$ referred by sum subsystem.
@@
If $X$ is an infinite set, and $k\in\mathbb{N}$, then for every finite coloring of $[X]^{k}$, there exists an infinite set $Y\subseteq X$ such that the set $[Y]^{k}$ is monochromatic.
The following theorem was due to K. Milliken [17] and A. Taylor [20] generalizing both Ramsey’s theorem and Hindman’s theorem, known as the Milliken-Taylor theorem.
@@
Let $m,r\in\mathbb{N}$, and $[\mathbb{N}]^{m}=\bigcup_{i=1}^{r}C_{i}$ be a $r$-coloring. Let $\langle x_{n}\rangle_{n}$ be a sequence in $\mathbb{N}$. Then there exists $i\in\{1,2,\ldots,r\}$, and a sum subsystem $FS(\langle y_{n}\rangle_{n})$ of $FS(\langle x_{n}\rangle_{n})$ such that $[FS(\langle y_{n}\rangle_{n})]_{<}^{m}\subseteq C_{i}.$

### v3

# Multidimensional Stronger Central Sets Theorem and its Polynomial Extension

@@ Abstract
We establish and fully characterize the multidimensional extension of the Stronger Central Sets Theorem, originally proved in [4]. Additionally, we develop a polynomial generalization of this result. Our approach utilizes tools from the algebra of the Stone–Čech compactification of discrete semigroups. Several applications of these results are also discussed.
@@
Keywords: Central sets, $C$-sets, Stronger Central Sets Theorem, Algebra of the Stone-Čech Compactifications of discrete semigroups.
@@ 1  Introduction
Throughout this article, we denote by $\mathbb{N}$ the set of all positive integers. For any $r\in\mathbb{N}$, an $r$-coloring of a set $X$ is a partition of $X$ into $r$ disjoint subsets. A coloring is said to be finite if it corresponds to an $r$-coloring for some $r\in\mathbb{N}$. A subset $F\subseteq X$ is called monochromatic if it is entirely contained within a single color class.
Let $(S,\cdot)$ be a semigroup, and let $\mathcal{F}$ be a family of subsets of $S$. The family $\mathcal{F}$ is said to be partition regular (P.R.) if, for any finite coloring of $S$, there exists a member $F\in\mathcal{F}$ that is monochromatic.
A fundamental result in arithmetic Ramsey theory is van der Waerden’s theorem [24], which states that the family of all arithmetic progressions of any given length is P.R. over $\mathbb{N}$. Another landmark result, due to Schur [21], asserts that the family
$\left\{\{x,y,x+y\}:x\neq y\in\mathbb{N}\right\}$
is P.R. over $\mathbb{N}$.
A nonlinear extension of van der Waerden’s theorem, known as the Polynomial van der Waerden Theorem, was established by Bergelson and Leibman in [3]. They employed techniques from topological dynamics and PET induction to prove this result. [^3] The authors actually proved a generalized form of the much stronger Szemerédi’s Theorem, but this aspect will not be discussed in the present paper.
Throughout this article, we denote by $\mathbb{P}$ the set of all polynomials with no constant term.
@@
Let $r\in\mathbb{N}$, and suppose that $\mathbb{N}$ is partitioned into $r$ color classes, i.e.,
$\mathbb{N}=\bigcup_{i=1}^{r}C_{i}.$
Then, for any finite collection of polynomials $F\subset\mathbb{P}$, there exist $a,d\in\mathbb{N}$ and some $1\leq j\leq r$ such that
$\{a+p(d):p\in F\}\subset C_{j}.$
An infinitary extension of Schur’s theorem motivates the following definition of IP sets.
Definition 1.2 (IP Set [14]).
  1. 1.
    For any set $X$, let $\mathcal{P}_{f}(X)$ denote the collection of all nonempty finite subsets of $X$.
  2. 2.
    Let $(S,+)$ be a commutative semigroup. A subset $A\subseteq S$ is called an IP set if there exists an injective sequence $\langle x_{n}\rangle_{n}$ in $S$ such that
    $A=FS\left(\langle x_{n}\rangle_{n}\right)=\left\{\sum_{n\in\alpha}x_{n}:\alpha\in\mathcal{P}_{f}(\mathbb{N})\right\}.$
    For brevity, given $\alpha\in\mathcal{P}_{f}(\mathbb{N})$, we write
    $x_{\alpha}=\sum_{n\in\alpha}x_{n}.$
A fundamental result in Ramsey theory, known as the Finite Sums Theorem or Hindman’s Theorem [9], asserts the following:
Theorem 1.3 (Hindman’s Theorem).
@@
For a comprehensive discussion on the classical approach to Ramsey theory, we refer the reader to [8]. In this article, we adopt an approach based on ultrafilter theory.
Before proceeding, we recall some fundamental results from the Stone–Čech compactification of discrete semigroups, a framework often referred to as topological algebra or the theory of ultrafilters.
A Brief Review of the Algebra of the Stone-Čech Compactifications of discrete semigroups
Ultrafilters are set-theoretic objects that are intimately related to Ramsey’s theory. In this section, we give a brief introduction to this theory. For details, we refer to the book of N. Hindman and D. Strauss [14] to the readers. A filter $\mathcal{F}$ over any nonempty set $X$ is a collection of subsets of $X$ such that
@@ A Brief Review of the Algebra of the Stone-Čech Compactifications of discrete semigroups
  • if $X=\bigcup_{i=1}^{r}A_{i}$ is any finite partition of $X$, then there exists $i\in\{1,2,\ldots,r\}$ such that $A_{i}\in p.$
Let $(S,\cdot)$ be any discrete semigroup. Let $\beta S$ be the collection of all ultrafilters. For every $A\subseteq S,$ define $\overline{A}=\{p\in\beta S:A\in p\}.$ Now one can check that the collection $\{\overline{A}:A\subseteq S\}$ forms a basis for a topology. This basis generates a topology over $\beta S.$ We can extend the operation $\cdot$" of $S$ over $\beta S$ as: for any $p,q\in\beta S,$ $A\in p\cdot q$ if and only if $\{x:x^{-1}A\in q\}\in p.$ With this operation $\cdot$", $(\beta S,\cdot)$ becomes a compact Hausdorff right topological semigroup. One can show that $\beta S$ is nothing but the Stone-Čech compactification of $S.$ Hence Ellis’s theorem guarantees that there exist idempotents in $(\beta S,\cdot)$. The set of all idempotents in $(\beta S,\cdot)$ is denoted by $E\left((\beta S,\cdot)\right).$ It can be shown that every member of the idempotents of $(\beta S,\cdot)$ contains an $IP$ set, which means every idempotent witnesses Hindman’s theorem. Using Zorn’s lemma one can show that $(\beta S,\cdot)$ contains minimal left ideals (minimal w.r.t. the inclusion). A well-known fact is that the union of such minimal left ideals is a minimal two-sided ideal, denoted by $K(\beta S,\cdot).$ Here we recall a few well-known classes of sets that are relevant to our work.
@@
  • AA is a thick set if for any finite subset $F\subset S$, there exists an element $x\in S$ such that $Fx=\{fx:f\in F\}\subset A$;
  • AA is a syndetic set if there exists a finite set $F\subset S$ such that $S=\bigcup_{x\in F}x^{-1}A$, where $x^{-1}A=\{y:xy\in A\}$;
  • AA is piecewise syndetic set if there exists a finite set $F\subset S$ such that $\bigcup_{x\in F}x^{-1}A$ is a thick set. It is well known that $A$ is piecewise syndetic if and only if there exists $p\in K(\beta S,\cdot)$ such that $A\in p.$
  • AA is central set if it belongs to a minimal idempotent in $\beta S$.
It can be proved that a set $A$ is thick if and only if there exists a left ideal $L$ such that $L\subseteq\overline{A}$. And a set $A$ is syndetic if and only if for every left ideal $L$, $L\cap\overline{A}\neq\emptyset.$ Note that if $f:(S,\cdot)\rightarrow(T,\cdot)$ be any map between discrete semigroups. Then $\tilde{f}:\beta S\rightarrow\beta T$ be the continuous extension of $f$ defined by $\tilde{f}(p)=\tilde{f}(\lim_{x\rightarrow p}x).$
@@
The tensor product of ultrafilters was introduced in [17]. In this section, we recall materials from [15]. Let $(S,\cdot)$ and $(T,\cdot)$ be two discrete semigroups.
@@
Using the induction argument, one can extend the above definition of tensor products for multiple ultrafilters.
@@
Similarly, we can readdress the above definition in the language of limit:
$\bigotimes_{i=1}^{k}p_{i}=\lim_{s_{1}\rightarrow p_{1}}\cdots\lim_{s_{k}\rightarrow p}(s_{1},\ldots,s_{k}).$
In [2], V. Bergelson, N. Hindman and K. Williams used the tensor product tricks to prove multi-dimensional Ramsey theoretic results including Milliken-Taylor theorem, and it’s polynomial extensions. For details on multi-dimensional Ramsey theoretic results we refer [2, 14, 18]. In [1], Beigelböck proved the simultaneous extension of the Central Sets Theorem and Milliken-Taylor theorem using a little different approach described in [2, Chapter 18]. One of our main motivations is to study the tensor product of minimal idempotent ultrafilters, which was not extensively covered earlier. As a consequence, we are able to prove the polynomial extensions of several multidimensional results in Ramsey theory. Before we proceed, we need the following technical Lemma from [2] (a variation can be found in [25]), which is essential for our purpose.
@@
[2, Lemma 2.5] Let $(S,\cdot)$ be a semigroup, let $m\in\mathbb{N}$, and let $p_{1},\ldots,p_{m}\in\beta S.$ Let $A\in\bigotimes_{j=1}^{m}p_{j}.$ Then for $j\in\{1,2,\ldots,m\}$ there exists $D_{j}:S^{j-1}\rightarrow\mathcal{P}(S)$ such that
@@
    for $j\in\{1,2,\ldots,m\},$ if for each $s\in\{1,2,\ldots,j-1\},$ $w_{s}\in D_{s}(w_{1},\ldots,w_{s-1}),$ then $D_{j}(w_{1},\ldots,w_{j-1})\in p_{j};$
@@
The Central Sets Theorem and its Polynomial Extensions
Before we proceed to our main results, we need to recall the Central Sets Theorem. After the foundation of both van der Waerden’s and Hindman’s theorem, an immediate question appeared if one can find a joint extension of both of these theorems. In [6], using the methods of Topological dynamics, Furstenberg defined the notions of Central Sets and proved that if $\mathbb{N}$ is finitely colored, then one of the color classes is Central. Then he proved a joint extension of both the van der Waerden’s and Hindman’s theorem, known as the Central Sets Theorem.
@@
[6, Proposition 8.21] Let $l\in\mathbb{N}$, and $A\subseteq\mathbb{N}$ be a central set. For each $i\in\{1,2,\ldots,l\}$ let $\langle x_{i,m}\rangle_{m=1}^{\infty}$ be a sequence in $\mathbb{\mathbb{Z}}$. Then there exist sequences $\langle b_{m}\rangle_{m=1}^{\infty}$ in $\mathbb{N}$ and $\langle K_{m}\rangle_{m=1}^{\infty}$ in $\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
@@
Before we state the stronger version of Theorem 1.8, we need the following definitions from [4].
@@
    AA is called a $J$-set if and only if for every $H\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, there exists $a\in\mathbb{N}$ and $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that for all $f\in H,$ $a+\sum_{t\in\beta}f(t)\in A.$
@@
It can be proved that $J$ sets are partition regular, and so $\mathcal{J}\neq\emptyset$. Again it is easy to verify that $(\mathcal{J},+)$ is a two sided ideal of $(\beta\mathbb{N},+)$, hence contains $K(\beta\mathbb{N},+).$
  • A set will be called $C$-set if it is a member of any idempotent in $(\mathcal{J},+)$.
As every piecewise syndetic set is a $J$ set, one can show that every Central set is a $C$ set. The following stronger version of the Central Sets Theorem was due to De, Hindman, and Strauss [4].
@@
Let $\tau=^{\mathbb{N}}\mathbb{N}$, and let $A\subseteq\mathbb{N}$.
@@
    AA is a $C$ set.
  2. 2.
    there exists functions $\alpha:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathbb{N},$ and $H:\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})\to\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
    1. (a)
@@
      $\sum_{i=1}^{m}\bigl(\alpha(G_{i})+\sum_{t\in H(G_{i})}f_{i}(t)\bigr)\in A.$
Polynomial Extension of the Central Sets Theorem
In [12, 13], Hindman and McCutcheon extensively studied the algebraic structure of VIP systems which are the generalizations of usual polynomials. They proved several results including the polynomial extension of Theorem 1.8. This was the joint extension of the polynomial van der Waerden’s theorem and the central sets theorem.
@@
[13] Let $F\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, let $T\in\mathcal{P}_{f}(\mathbb{P})$ and let $A$ be a central subset of $\mathbb{N}$. Then there exist sequences $\langle b_{n}\rangle_{n=1}^{\infty}$ in $\mathbb{N}$ and $\langle H_{n}\rangle_{n=1}^{\infty}$ in $\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that
@@
Later in [7], the authors found the polynomial extension of Theorem 1.10, which is the stronger version of Theorem 1.11. To state this result we need the following definition which polynomialize the notion of $J$ sets.
@@
    AA is called a $J_{p}$-set if and only if for every $F\in\mathcal{P}_{f}(\mathbb{P})$, and every $H\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, there exists $a\in\mathbb{N}$ and $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ such that for all $P\in F$ and all $f\in H,$ $a+P\left(\sum_{t\in\beta}f(t)\right)\in A$
@@
    if $n\in\mathbb{N}$, $G_{1},G_{2},\cdots,G_{n}\in\mathcal{P}_{f}(^{\mathbb{N}}\mathbb{N})$, $G_{1}\subsetneq G_{2}\subsetneq\cdots\subsetneq G_{n}$ and for all $i\in\{1,2,\ldots,n\}$, $f_{i}\in G_{i}$, then for all $P\in T$,
@@
However likewise in Theorem 1.10, the converse direction of Theorem 1.13 is not known. The proof of Theorem 1.13 uses induction argument. By observing that any $\beta\in\mathcal{P}_{f}\left(\mathbb{N}\right)$ is a subset of $\{1,\dots,n\}$ for some large enough $n\in\mathbb{N}$, by going along the lines of the proof of Theorem 1.13, one can deduce the following seemingly stronger version of Theorem 1.13.
@@
Milliken-Taylor Theorem and its Twofold Extensions
In 19751975-$76$, Milliken and Taylor independently proved the joint extension of both Hindman theorem and Ramsey’s theorem. To state the Milliken-Taylor Theorem, we need to introduce some notation. Given $F,G\in\mathcal{P}_{f}(\mathbb{N})$, we write $F<G$ to mean that $\max F<\min G.$ When we say that a sequence $\langle H_{n}\rangle_{n}$ in $\mathcal{P}_{f}(\mathbb{N})$ is increasing if for every $n\in\mathbb{N}$, $\max H_{n}<\min H_{n+1}.$ We present the different terminology because these special cases arise frequently.
@@
    $FS(\langle x_{n}\rangle_{n})]_{<}^{k}=\left\{\left(\sum_{t\in F_{1}}x_{t},\sum_{t\in F_{2}}x_{t},\ldots,\sum_{t\in F_{k}}x_{t}\right):F_{1}<F_{2}<\cdots<F_{k}\right\}.$
@@
    In a semigroup $(S,\cdot),$ $FP(\langle y_{n}\rangle_{n})$ is a product subsystem of $FP(\langle x_{n}\rangle_{n})$ if and only if there exists an increasing sequence $\langle H_{n}\rangle_{n}$ in $\mathcal{P}_{f}(\mathbb{N})$ such that for each $n\in\mathbb{N},$ $y_{n}=\Pi_{t\in H_{n}}x_{t}$ where the products $\Pi_{t\in H_{n}}x_{t}$ are computed in increasing order of indices. If the semigroup is commutative, we denote it by $(S,+)$ and the product subsystem $FP(\langle y_{n}\rangle_{n})$ is denoted by $FS(\langle y_{n}\rangle_{n})$ referred by sum subsystem.
@@
If $X$ is an infinite set, and $k\in\mathbb{N}$, then for every finite coloring of $X]^{k}$, there exists an infinite set $Y\subseteq X$ such that the set $Y]^{k}$ is monochromatic.
The following theorem was due to K. Milliken [19] and A. Taylor [23] generalizes both Ramsey’s theorem and Hindman’s theorem, known as the Milliken-Taylor theorem. They originally proved two different but equivallent versions of the Milliken-Taylor theorem (see [2]). Here we quote the version from [19].
@@
Let $m,r\in\mathbb{N}$, and $\mathbb{N}]^{m}=\bigcup_{i=1}^{r}C_{i}$ be a $r$-coloring. Let $\langle x_{n}\rangle_{n}$ be a sequence in $\mathbb{N}$. Then there exists $i\in\{1,2,\ldots,r\}$, and a sum subsystem $FS(\langle y_{n}\rangle_{n})$ of $FS(\langle x_{n}\rangle_{n})$ such that $FS(\langle y_{n}\rangle_{n})]_{<}^{m}\subseteq C_{i}.$